# Rethinking Diffusion for Text-Driven Human Motion Generation

### v1

@@ Appendix A Detailed Deduction
Appendix B Implementation Details
For our method, the AutoEncoder is a 3-layer ResNet-based encoder-decoder with a hidden dimension of 512 and a total downsampling rate of 4. For the generation branch, we utilize a single-layer AdaLN-Zero transformer encoder with a hidden dimension of 1024 and 16 heads as our masked autoregressive transformer. The diffusion MLPs consist of 16 layers with a hidden dimension of 1792. We also present the model scalability results in Sec. C.2.
@@ Appendix B Implementation Details
Appendix C Additional Quantitative Results
C.1 AutoEncoder Reconstruction Results
@@
C.2 Model Scalability
We train six versions of our proposed model (DDPM approach), varying three transformer sizes and four diffusion MLP sizes (S, M, L, XL). These models range in size from around 30M, 100M, 180M, to 290M parameters. The performance results are summarized in Sec. A.1. We observe that increasing the model size, particularly the diffusion MLPs size, improves overall generation performance, especially in terms of FID.
Appendix D Temporal Editing
Our method is capable of performing temporal editing in a zero-shot manner (i.e. utilizing the model trained for text-to-motion generation without any editing-specific fine-tuning). In our method, temporal motion editing is easily achieved by treating the latents that need to be edited as masked latents and then generating motions following our standard generation procedure in Sec. 3.2 which is conditions on the unmasked tokens (i.e. non-edit latents) and the editing textual instructions. We visually illustrate this process in Fig. A1 and we also include temporal editing results in the locally-run, anonymous HTML file referenced in Appendix E.
Appendix E Additional Qualitative Results

### v2

@@ Appendix B Detailed Related Works
Table A2: Further Ablation Study and Optimization Routine.
[…]
Top-1 Top-2 Top-3
MDM [88]-50Step-$\boldsymbol{\epsilon}$ $31.265$ $0.054$ $0.103$ $0.147$
+Masked AR $2.196$ $0.387$ $0.595$ $0.703$
++Essential Only $0.657$ $0.475$ $0.668$ $0.774$
$+$AE (Ours) $\mathbf{0.116}$ $\mathbf{0.492}$ $\mathbf{0.690}$ $\mathbf{0.790}$
++++$\mathbf{X}_{0}$-Pred $0.135$ $0.485$ $0.686$ $0.784$
++++Velocity (Ours) $\mathbf{0.114}$ $\mathbf{0.500}$ $\mathbf{0.695}$ $\mathbf{0.795}$
Table A3: Training Baseline Methods with Reformed Motion Data Representation and Distribution, Linear schedule and $\boldsymbol{\epsilon}$-prediction
[…]
Top-1 Top-2 Top-3
MDM [88] $1.574$ $0.279$ $0.336$ $0.415$
MDM [88]-Essential $0.753$ $0.436$ $0.627$ $0.732$
MotionDiffuse [113] $0.778$ $0.450$ $0.641$ $0.753$
MotionDiffuse [113]-Essential $0.533$ $0.459$ $0.650$ $0.757$
MDM [88]-Latent $0.327$ $0.475$ $0.663$ $0.768$
Table A4: Original Evaluator Results on HumanML3D.
[…]
Top-1 Top-2 Top-3
GT $0.002$ $0.511$ $0.703$ $0.797$
[…]
MDM [88]-50Step $0.489$ $0.455$ $0.645$ $0.749$
MDM [88]-50Step-Reproduce $0.481$ $0.459$ $0.651$ $0.753$
T2M-GPT [111] $0.141$ $0.492$ $0.679$ $0.775$
T2M-GPT [111]-Reproduce $0.115$ $0.497$ $0.685$ $0.779$
MMM [76] $0.089$ $0.515$ $0.708$ $0.804$
MMM [76]-Reproduce $0.071$ $0.517$ $0.711$ $0.805$
MoMask [26] $\mathbf{0.045}$ $\underline{0.521}$ $\underline{0.713}$ $\underline{0.807}$
MoMask [26]-Reproduce $0.093$ $0.508$ $0.701$ $0.796$
Ours $\underline{0.061}$ $\mathbf{0.523}$ $\mathbf{0.715}$ $\mathbf{0.810}$
@@
Appendix C Implementation Details
For our method, the AutoEncoder is a 3-layer ResNet-based encoder-decoder with a hidden dimension of 512 and a total downsampling rate of 4. For the generation branch, we utilize a single-layer AdaLN-Zero transformer encoder with a hidden dimension of 1024 and 16 heads as our masked autoregressive transformer. The diffusion MLPs consist of 16 layers with a hidden dimension of 1792. We also present the model scalability results in Sec. D.5.
@@ Appendix C Implementation Details
Appendix D Additional Quantitative Results
D.1 AutoEncoder Reconstruction Results
@@
D.2 Baseline Methods Training With Reformed Data Representation and Distribution
In Tab. A3, we demonstrate that training baseline methods using only essential dimensions can already lead to significant improvements, and processing into latent space may further enhance results.
D.3 Original Evaluation Results
The original evaluator is flawed due to the unnecessary focus on redundant motion representations and the new evaluators are proposed to deal with this issue. Therefore, we strongly discourage utilizing the original evaluation method to access all methods. Also, using the original evaluator requires additional processing to convert our outputs to joints and back to the redundant representations. This inevitably introduces errors, and loses one motion frame (from joints to HumanML3D representations), and thus unfairly penalizes our method. Nevertheless, we include results in Tab. A4. Note, our method is converted and thus may be biased and inaccurately evaluated under this unfair circumstance. And notably, our reproduced MoMask exhibits worse results, similar to issues reported in their GitHub (Issues 27, 43, 99), and even ground truth motions were penalized due to the additional operations.
D.4 Further Ablation Study and Optimize Routine
In Tab. A2, we provide a further ablation study and an optimization routine starting from an MDM-based cosine schedule, $\boldsymbol{\epsilon}$ prediction approach to our approach. The results demonstrate the advantage of masked regression over original diffusion and the importance of our further optimization (motion representation reformation) over pure adoption of image MAR.
D.5 Model Scalability
We train six versions of our proposed model (DDPM approach), varying three transformer sizes and four diffusion MLP sizes (S, M, L, XL). These models range in size from around 30M, 100M, 180M, to 290M parameters. The performance results are summarized in Appendix B. We observe that increasing the model size, particularly the diffusion MLPs size, improves overall generation performance, especially in terms of FID.
Appendix E Temporal Editing
Our method is capable of performing temporal editing in a zero-shot manner (i.e. utilizing the model trained for text-to-motion generation without any editing-specific fine-tuning). In our method, temporal motion editing is easily achieved by treating the latents that need to be edited as masked latents and then generating motions following our standard generation procedure in Sec. 3.2 which is conditions on the unmasked tokens (i.e. non-edit latents) and the editing textual instructions. We visually illustrate this process in Fig. A1 and we also include temporal editing results in the locally-run, anonymous HTML file referenced in Appendix F.
Appendix F Additional Qualitative Results